Chapter 33, Problem 023


Someone plans to float a small, totally absorbing sphere 0.514 m above an isotropic point source of light, so that the upward radiation force from the light matches the downward gravitational force on the sphere. The sphere's density is 21.4 g/cm3, and its radius is 2.08 mm. (a) What power would be required of the light source? (b) Even if such a source were made, would the support of the sphere be unstable? Assume that the speed of light is 2.998*108 m/s.
Answer #1

answer) first we have the condition

W=Fr........................1)

we have intensity I=P/Ar and Ps=I/c

so substituting the value of I

Ps=P/cAr..............2)

now we have the pressure on the sphere Ps=Fs/A

so Fs=PsA and from eqn 2 we have

Fs=PAs/cAr...............3)

we know the relation

W=mg

so eqn 3 becomes

mg=PAs/cAr

P=mgArc/As=\rhoVsg4pid2/pir2=\rho*(4/3)pir3g4pid2c/pir2=21400*(4/3)pi*(0.00208)3*9.8*4pi*0.5142*2.998*108/pi*(0.00208)2=5.78*1011W

so answer is 5.78*1011W or 5.8*1011 or 6*1011W.

b) yes the support will be unstable because the equal and opposite forces which will be keep the sphere in equilibrium will not be on the same line.